OpenStudy (anonymous):

The surface area of an oil spill gets 105% larger every day, represented by the function s(x) = (1.05)x − 1. On the first day, it covered an area of 31 square meters. Which function would be used to find the area of the oil spill on the 15th day

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well there seems to be something missing in the way you have written the question. is it \[s(x) = 31 \times\ (1.05)^{x -1}\]

OpenStudy (anonymous):

yes @campbell_st

OpenStudy (campbell_st):

ok... so x = 15, just substitute and calculate to find the surface area

OpenStudy (anonymous):

I know but my choices are: s(x)=1.05(31)^x-1 or s(x)=31(1.05)^x-1 would it be the second choice?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

its the 2nd choice